OpenStudy (anonymous):

What is the solution of the following system? -3x-2y=-12 9x+6y=-9

OpenStudy (anonymous):

1.(2,1) 2.no solutions 3.(-2,-1) 4.infinitely many solutions

OpenStudy (anonymous):

what do you want us to find?

OpenStudy (anonymous):

ohhh, hold on...

OpenStudy (anonymous):

I would say D, but check with @johnweldon1993 hes amazing;)

OpenStudy (anonymous):

@reff23

OpenStudy (anonymous):

I was thinking that too because they are parallel

OpenStudy (johnweldon1993):

-3x-2y=-12 9x+6y=-9 So we can make the top equation look like -2y = 3x - 12 y = (-3/2)x + 6 and the second would be 6y = -9x - 9 y = (-3/2)x - 3/2 Now you were correct in saying these lines are parallel However the solution to a system of equations is where the 2 lines intersect... do parallel lines (providing they aren't the same line) ever intersect?

OpenStudy (anonymous):

@reff23 hes trying to help you

OpenStudy (anonymous):

No, I don't think they ever intersect.

OpenStudy (johnweldon1993):

^Correct, so the correct choice here will be B, there will be NO point where 2 parallel lines will intersect

OpenStudy (anonymous):

Thank you.

OpenStudy (johnweldon1993):

For the case of choice D that would be correct if the 2 equations you got are the same equation for example 2x + 4y = 6 4x + 8y = 12 If you simplify those down, you get the same line... hence they would intersect at EVERY point, or in other words, have infinitely many solutions